Which of the following pairs represent the same rational number? 
(-16/20) and (20/-25)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We have to check the given pair represents the same rational number.
Then,
-16/20 = 20/-25
-4/5 = 4/-5
∵-4/5 = -4/5
∴-16/20 = 20/-25
So, the given pair is represents the same rational number.

Thanks